limh→ 0 (6+h)


2−36


h

1
Expert's answer
2022-05-02T15:41:20-0400
"\lim\limits_{h\to 0}\dfrac{(6+h)^2-36}{h}=\lim\limits_{h\to 0}\dfrac{36+12h+h^2-36}{h}"

"=\lim\limits_{h\to 0}(12+h)=12+0=12"